### v3

Abstract
Forecast stability, that is, the consistency of predictions over time, is essential in business settings where sudden shifts in forecasts can disrupt planning and erode trust in predictive systems. Despite its importance, stability is often overlooked in favor of accuracy. In this study, we evaluate the stability of point and probabilistic forecasts across several retraining scenarios using three large forecasting datasets and ten different global forecasting models. To analyze stability in the probabilistic setting, we propose a new model-agnostic, distribution-free, and scale-free metric that measures probabilistic stability: the Scaled Multi-Quantile Change (SMQC). The results show that less frequent retraining not only preserves but often improves forecast stability, challenging the need for frequent retraining. Moreover, the study shows that accuracy and stability are not necessarily conflicting objectives when adopting a global modeling approach. The study promotes a shift toward stability-aware forecasting practices, proposing a new metric to evaluate forecast stability effectively in probabilistic settings, and offering practical guidelines for building more stable and sustainable forecasting systems.
keywords:
Time series, Forecasting competitions, Cross-learning, Global models, Forecast stability , Vertical stability , Machine learning, Deep learning, Conformal predictions
journal: International Journal of Forecasting
1 Introduction
In recent years, global forecasting models, those trained across multiple time series simultaneously, have emerged as a powerful alternative to traditional local approaches, particularly in large-scale applications such as retail demand and energy forecasting. While their accuracy and efficiency have been widely studied, less attention has been paid to the stability of their forecasts over time. In many operational contexts, forecasts are produced regularly and serve as the foundation for critical business decisions, from inventory planning to resource allocation. In such settings, forecast stability becomes a key requirement; not only should forecasts be accurate, but they should also remain reasonably stable over time. Indeed, forecast stability refers to the consistency of predictions produced by a forecasting model as new data becomes available. As defined by Godahewa et al. (2025), forecast stability can be categorized into two primary forms: vertical stability, which concerns the consistency of forecasts for the same target date across different forecast origins, and horizontal stability, which addresses the smoothness of forecasts across the forecast horizon from a single origin. Both types play a distinct role: vertical stability helps avoid costly forecast revisions that can disrupt planning cycles, while horizontal stability prevents erratic behavior across time steps that could lead to inefficient operations and amplification of demand fluctuations, such as the bullwhip effect in supply chains (Lee et al., 1997). In particular, vertical stability may also be seen as temporal robustness, in the sense that a forecasting model is stable to updates (or retraining) as new observations are available, a property that is critical in practice for avoiding frequent changes in business decisions. Forecast instability, indeed, can have serious consequences. Unstable forecasts may lead to frequent and expensive adjustments in supply chain plans, diminish trust in the forecasting system, and complicating the decision-making processes, leading to suboptimal business outcomes.
Nevertheless, despite the importance of forecast stability, it is still common practice to evaluate forecasting models only on their accuracy, even in frequent retraining settings where stability is more likely to be compromised. This can also be due to an inherent perception that there exists a trade-off between accuracy and stability. As new information becomes available, updated forecasts may naturally differ from earlier versions, ideally improving in accuracy at the expense of stability. Conversely, overly stable forecasts might sacrifice accuracy by ignoring new, valuable information. Ensembling techniques, which integrate multiple forecasts from different models, have been proposed as a mechanism to concurrently reduce bias and variance (Wang et al., 2023). By aggregating diverse predictions, ensembling methods can alleviate the impact of individual model volatility and potentially achieve a more favorable equilibrium between accuracy and stability, but at the expense of a severe increase in computational costs. Even if the trade-off between forecasting accuracy and stability presents a significant challenge for organizations seeking to balance predictive performance with operational stability, recent research suggests that high levels of stability and accuracy can be achieved at the same time ((Godahewa et al., 2025), (Van Belle et al., 2023)). Moreover, Zanotti (2025) demonstrated that the retraining frequency does not harm the forecast accuracy of global models, suggesting that, when the data-generating process remains relatively stable, global models can preserve predictive performance even when updated less frequently. This finding raises a natural question regarding forecast stability. Since retraining inherently produces new model artifacts, the frequency with which models are retrained may influence the temporal consistency of forecasts. Whether less frequent retraining systematically improves, worsens, or leaves unchanged the stability of global forecasting models is therefore an important question that needs to be addressed.
1.1 Research Question
We aim to address the question ”How does the retraining frequency of global forecasting models affect the stability of their forecasts?”. Purposely, we study whether the global modeling approach produces stable forecasts, and we try to understand the effects of retraining on the forecast stability, that is, whether avoiding re-estimation for every new observation damages the stability of global models. To address this question, we rely on three of the most recent and comprehensive forecasting datasets: the M4, the M5, and the VN1 competition data.
To generally understand the stability of the forecasting models, we consider ten distinct global forecasting methods (five from the traditional machine learning domain and five based on commonly used deep neural network architectures). We also compared the results with two well-known univariate local forecasting models. Moreover, we examine stability across a range of retraining scenarios, continuous retraining to no retraining, by exploring periodic strategies that broadly encompass the most practical and effective approaches.
1.2 Contributions
Our contribution is fourfold:
  • 1.
    We provide the first comprehensive study of the forecast stability of global models, using 10 distinct methods, a diverse collection of real-world datasets, and evaluating both point and probabilistic predictions.
  • 2.
    We analyze the relationship between the retraining period and forecast stability, comparing different scenarios to quantify the impact of frequent retraining in terms of the stability of forecasting.
  • 3.
    We suggest new metrics to evaluate the stability of probabilistic forecasts.
  • 4.
    We present practical guidelines for organizations and practitioners on when and how often to retrain global forecasting models to obtain stable forecasts.
By tackling these aspects, this paper contributes to both the forecasting and machine learning communities by providing insights into the stability of global forecasting models.
1.3 Overview
The rest of this paper is organized as follows. After a brief review of related works (Section 2), in Section 3 we describe the design of the experiment used in our study. The datasets and their characteristics are presented in 3.1, and the methods adopted for global forecasting are discussed in 3.2. The concepts related to model update and retrain scenario are explained in 3.3, together with the evaluation strategy adopted, while the metrics used to assess the stability of models are shown in 3.4. In Section 4 we discuss the empirical findings of our study, including forecast accuracy and stability on the different scenarios. Finally, Section 5 contains our summary and conclusions.
2 Related works
2.1 Global modeling
The cross-learning approach has seen substantial development in recent years. Today, most time series forecasting studies include at least a benchmark comparison involving global models (GMs), underscoring their growing importance in the field.
We define $\mathcal{Y}$ as the set of all available time series in a dataset, where each $Y_{i}\in\mathcal{Y}$ represents an individual time series. Let $\mathcal{F}$ denote the set of possible predictive functions, with each $F\in\mathcal{F}$ corresponding to a specific forecasting model. Without loss of generality, we assume that all necessary information for prediction is contained within $\mathcal{Y}$. Under the local modeling paradigm (LMs), forecasts for a horizon $h$ are generated by training a separate model for each individual time series. This implies that each series $Y_{i}$ is associated with its own model, characterized by its own set of parameter values.
$Y_{i}^{h}=F(Y_{i},\theta_{i})\mbox{.}$ (1)
In contrast, under the global modeling framework, forecasts for each individual time series are generated using a single model trained on the entire dataset. This approach leverages cross-series information, allowing the model to learn shared patterns and structures across all time series in $\mathcal{Y}$.
$Y_{i}^{h}=F(\mathcal{Y},\Theta)\mbox{.}$ (2)
It is important to note that in the cross-learning methodology, the model parameters $\Theta$ are not specific to individual time series but are shared across all series in the dataset. This parameter sharing is a key characteristic of global models, enabling them to generalize patterns across series and potentially reducing the forecast instability.
Semenoglou et al. (2021) demonstrated the high accuracy of GMs on the M4 competition dataset, while Hewamalage et al. (2022) explored the conditions under which global models are competitive. Additionally, Montero-Manso and Hyndman (2021) and (Montero-Manso, 2023) provided theoretical support showing that GMs can match or surpass the accuracy of LMs, with lower complexity and without assuming data similarity. Global models have proven to be the most accurate method in several forecasting domains, including retail demand (Spiliotis et al. (2022), Bandara et al. (2019), Juan R Trapero and Fildes (2015)), electricity demand (Buonanno et al., 2022), water demand (Groß and Hans, 2024), gas consumption (Gaweł and Paliński, 2024), and crop production (Ibañez and Monterola, 2023). Their effectiveness was particularly evident during the M5 competition (Makridakis et al., 2022b), where tree-based models employing cross-learning ranked among the top-performing solutions (Januschowski et al., 2022). To further enhance GM performance, several strategies such as clustering (Godahewa et al. (2021a), Bandara et al. (2020)) and data augmentation (Bandara et al., 2021) have been explored. Moreover, new machine learning (Godahewa et al., 2023) and deep learning (Oreshkin et al., 2020) architectures have been specifically designed to support cross-learning. Recently, research has begun focusing on improving GMs’ ability to capture local patterns (Sen et al., 2019) and enhance their explainability (Rajapaksha et al., 2023).
2.2 Forecast stability
From a forecasting stability perspective, the literature is impoverished. Godahewa et al. (2025) introduced the first classification of the different types of forecasting stability, that is horizontal or vertical stability, proposing a new model-agnostic framework based on linear combinations of predictions to obtain more stable forecasts. Van Belle et al. (2023), instead, extended an existing deep learning architecture (NBEATS) to optimize forecasts from both a traditional forecast accuracy perspective, as well as a forecast stability perspective, directly including an instability component into the loss function of the model. There is also active research on forecast stability within judgmental forecasting ground (Fildes and Goodwin, 2021). However, almost all the literature related to forecast stability is focused on point prediction stability only. This is not due to the absence of probabilistic distance measures, nor to the lack of probabilistic forecasting techniques (Makridakis et al., 2022a). Rather, probabilistic stability has received limited attention because existing evaluation metrics are typically designed to compare forecasts to realizations or to assess calibration and sharpness, rather than to quantify temporal consistency across successive forecast origins. Nevertheless, in many forecasting applications, such as supply chain management, it is crucial to generate and evaluate predictions probabilistically, whether through prediction intervals, quantiles, or full predictive distributions (Fildes et al., 2022). The seminal work of Van Belle et al. (2024) is the only one addressing the problem of probabilistic stability evaluation, but it only applies to a specific forecasting model (NBEATS), which has been modified to output a full predictive Gaussian distribution. Therefore, the stability evaluation of probabilistic forecasts is an open question, especially in a model-agnotic way. Luckily, among the methods developed for uncertainty quantification, Vovk et al. (2005) introduced Conformal Inference, which is a model-agnostic framework that offers valid uncertainty estimates and can also be applied in time series forecasting settings (Stankeviciute et al., 2021).
2.3 Model retraining
In the context of model retraining and updating strategies, the most comprehensive work in time series forecasting is by Zanotti (2025), who showed the positive effects of reducing retraining of global models on both accuracy and forecasting costs. The seminal work of Spiliotis and Petropoulos (2024) extensively investigated the impact of various retraining scenarios and parameter update methods on model performance, focusing on the exponential smoothing family of models, following the traditional local modeling approach. (Huber and Stuckenschmidt, 2020) touched on retraining within the retail demand domain, but the study was restricted to a small set of models, limited retraining strategies, and a proprietary daily dataset. While these findings are encouraging, there has been little direct investigation into whether global models specifically can retain stability with less frequent updates.
Building on these prior studies, our work directly examines the stability of global models. Moreover, by systematically evaluating a wide range of retraining strategies and their effects on the forecasting stability of various global models, we seek to offer both theoretical insights and practical guidance for promoting more stable forecasting practices.
3 Experimental design
This section presents the empirical analysis carried out to investigate the stability of global models and to determine whether less frequent retraining scenarios can yield stability outcomes comparable to those of the baseline scenario, which involves the most frequent retraining. We begin by describing the datasets used in our experiments, followed by an overview of the machine learning and deep learning models employed. Finally, we detail the instability measures, the various retraining scenarios considered, and the evaluation strategy applied to assess forecast performance.
3.1 Datasets
For our experiments, we employed three forecasting datasets: the M4, the M5, and the VN1 competition datasets. The M4 competition is part of the well-known M-competitions series organized in 2018 by Spyros Makridakis and colleagues Petropoulos and Makridakis (2020). It includes real-world time series drawn from diverse domains such as finance, economics, demographics, industry, and other indicators, and spans multiple frequencies. In our study, we focused on the 4,227 daily time series characterized by substantial variability in length, seasonality, and noise structure. The heterogeneous nature of the series makes it particularly well-suited for studying general retraining behaviors that are not domain-specific. The M5 competition, also part of the M-competitions series, was organized to benchmark forecasting methods in the context of retail demand (Makridakis et al., 2022c). The M5 dataset (Howard and Makridakis, 2020) includes 30,490 daily time series representing unit sales of Walmart products across three categories (Food, Hobbies, and Household) sold in ten stores located in California, Texas, and Wisconsin. The data span from 2011 to 2016 and are characterized by high intermittency and a hierarchical structure, enabling forecasts at multiple aggregation levels (e.g., SKU, product category, store, and state). The dataset also includes exogenous variables such as prices, promotions, and special events (e.g., holidays), which can influence demand. The VN1 Forecasting - Accuracy Challenge, organized in October 2024 by Flieber, Syrup Tech, and SupChains, represents the first edition of a new competition series (Vandeput, 2024). The dataset comprises weekly sales data for 15,053 products sold by U.S.-based e-vendors between 2020 and 2024. Unlike the M5, which involves a single retailer (Walmart) and a limited number of physical stores, the VN1 dataset captures sales from 328 warehouses operated by 46 different retailers. As far as we are aware, our study is among the first to evaluate forecasting models on this dataset. Together, the M5 and the VN1 datasets offer the most recent and comprehensive time series collections related to retail demand, enabling good generalizability of our findings also to such domain.
3.2 Forecasting models
In our study, we focused on analyzing the stability of global methods, as our primary goal is to assess whether this modeling approach, unlike the traditional local one, can maintain performance with less frequent retraining. Nevertheless, we also compared the results with two popular local models, that is the Exponential Smoothing (ETS) (Hyndman et al., 2002) and the Autoregressive Integrated Moving Average (ARIMA) (Hyndman and Khandakar, 2008) methods in their automatic versions  [^1] A comprehensive analysis of retraining effects on the stability of local models was out of the scope of this study. . To conduct a comprehensive evaluation of the stability of global models, we incorporated both traditional machine learning algorithms and state-of-the-art deep learning architectures. The selected models are well-established in the time series forecasting literature and represent a range of methodological paradigms, allowing for a broad and informative comparison.
Machine learning models have proven effective in forecasting tasks due to their ability to capture complex non-linear relationships in data. They are also relatively easy to train and tend to deliver strong performance when leveraging cross-learning techniques. In this study, we experimented with Linear (Pooled) Regression and four widely used tree-based methods. While machine learning models are easier to train compared to deep learning alternatives, they often rely heavily on high-quality feature engineering (Januschowski et al., 2022). For our experiments, we adopted simplified versions of the feature engineering pipelines used in the top-performing M4, M5 and VN1 solutions. We constructed time series features such as lags, rolling means, expanding means, and calendar features (e.g., year, month, week, day of the week). For the M5 and VN1 datasets, we also included static metadata such as store, product, category, and location identifiers. For the M5 dataset, additional external features like special events were also incorporated. Model hyperparameters were selected based on configurations used in top-performing competition solutions when available; otherwise, we relied on the recommended defaults from the respective automatic libraries. We compared five machine learning models for time series forecasting. Linear Regression (LR) is a classical statistical model effective with appropriate feature engineering, often used as a benchmark in global model performance ((Montero-Manso and Hyndman, 2021), (Godahewa et al., 2021b)). Random Forest (RF), an ensemble learning method, captures non-linear patterns and is robust to overfitting, making it suitable for demand forecasting ((Breiman, 2001), (Januschowski et al., 2022)). Extreme Gradient Boosting (XGBoost) and Light Gradient Boosting Machine (LGBM) are gradient-boosted models known for speed and accuracy, with LGBM excelling in large, high-dimensional datasets ((Chen and Guestrin, 2016), (Ke et al., 2017), Makridakis et al. (2022b)). Categorical Boosting (CatBoost) specializes in handling categorical data with minimal preprocessing (Prokhorenkova et al., 2018).
Deep learning models have gained significant traction in time series forecasting due to their ability to model long-range temporal dependencies and to learn hierarchical representations directly from raw input data (Goodfellow et al., 2016). Unlike machine learning models, deep learning architectures typically do not require extensive manual feature engineering. They can autonomously learn lagged structures, trend, and seasonality directly from the raw series. Nevertheless, training deep models is often more complex due to the higher number of hyperparameters and their sensitivity to configuration choices, which can significantly affect performance (Smyl, 2020). In our implementation, we followed competition-proven practices by providing only static, calendar, and external covariates, together with historical values, as inputs. Top solutions’ guidelines were adopted for setting the model hyperparameters when available, and otherwise, the recommended defaults from the respective automatic libraries were used. We compared five deep learning models for time series forecasting. Multi-Layer Perceptron (MLP) is a versatile, efficient neural network used in many time series forecasting tasks (Rosenblatt, 1958), while Recurrent Neural Networks (RNNs), particularly LSTMs and GRUs, capture temporal dependencies in sequential data ((Cho et al., 2014), (Hochreiter and Schmidhuber, 1997)). Temporal Convolutional Networks (TCN) offer an alternative by using causal convolutions to capture long-range dependencies (Van den Oord et al., 2016). Neural Basis Expansion Analysis for Time Series (NBEATS) and Neural Hierarchical Interpolation for Time Series (NHITS) are state-of-the-art models for time series forecasting, with NBEATS offering interpretable trend and seasonality components (Oreshkin et al., 2020), and NHITS improving upon NBEATS with hierarchical interpolation mechanisms (Challu et al., 2022).
All models were implemented in Python using Nixtla’s framework (Nixtla, 2022). Specifically, the statsforecast library was adopted to train univariate statistical models, the mlforecast library was used to train the global machine learning models, while neuralforecast was employed for efficiently training the global deep learning models.
For this study, we adopted a cloud computing machine NC6s_v3 hosted on Microsoft Azure, with Linux Ubuntu 24 operating system, 1 Graphical Processing Unit, 6 Computing Processing Units, 112GB of memory.
3.3 Evaluation strategy
Out-of-sample evaluation is a cornerstone of time series forecasting, providing a way to test a model’s generalization ability beyond the training period, an essential step given that future data may diverge from historical patterns due to structural breaks, level shifts, or unexpected shocks (Tashman, 2000). Among the various evaluation strategies, the rolling origin evaluation has emerged as the most widely accepted and rigorous method (Bergmeir and Benítez, 2012). It respects the chronological nature of time series data while enabling repeated assessments across multiple forecast origins. In this framework, the time series is split into a training set and a test set, where the model is trained on the former and evaluated on the latter. Forecasts are generated for a defined horizon $h$, and at each iteration, the forecast origin is shifted forward by a specified step size, typically one, to simulate real-time forecasting. The model is then retrained on the updated training data, using either a fixed-length or expanding window. Evaluation metrics (see Section 3.4) are averaged over all iterations to provide a robust estimate of forecasting accuracy.
Compared to fixed origin evaluations, rolling origin offers a more nuanced view of a model’s robustness by exposing it to a range of temporal conditions (including seasonal patterns, trends, and potential anomalies) that may not be captured in a single evaluation window (Bergmeir and Benítez, 2012). This iterative approach reduces the risk of overfitting to one specific train-test split, and is especially beneficial in operational contexts like retail, energy, logistics, or finance, where forecasts are regularly updated in response to incoming data. It also aligns well with real-world decision-making processes that require forecasts to adapt over time.
In practice, the design of a rolling origin evaluation depends on several parameters: the size of the test set, the length of the forecast horizon, the step size, and the windowing strategy (fixed vs expanding). The expanding window strategy is particularly suited for short time series, as it maximizes the amount of historical data available at each iteration. This approach is widely used in applied settings (Petropoulos and et al., 2022), and it is the one we adopted in our study, especially given the relatively short length of the weekly VN1 series. The Table 1 outlines the key parameters used in our experiments, including a test set spanning at least one full year, to mitigate seasonal bias, horizons aligned with typical business use cases, and a step size of one to maximize evaluation frequency. A step size of one is also relevant in the context of stability evaluation, since multiple forecasts for the same target are required to compute stability metrics.
Table 1: Retraining set, train and test size, frequency, forecast horizon, and step size for each dataset. The setup is based on the respective time series frequency.
M4 / M5 VN1
Frequency ($f$) daily (7) weekly (52)
Train size ($n$) $\geq 730$ $\geq 157$
Test size ($T$) 364 52
Horizon ($h$) 28 13
Step size ($p$) 1 1
Retrain set ($r$) {7, 14, 21, 30, 60, 90, 120, 150, 180, 364} {1, 2, 3, 4, 6, 8, 10, 13, 26, 52}
Furthermore, to investigate the impact of retraining on forecasting stability, we examined various retraining scenarios, or retraining windows. As defined in Zanotti (2025), each scenario $r$ indicates the number of new observations after which the model is retrained. These retraining scenarios are tailored to the data frequency, daily for M4 and M5, and weekly for VN1, since frequency dictates both the forecast horizon and business review cycles. For instance, in the daily case, $r=7$ reflects weekly updates. Across all scenarios, we trained global models on each dataset, updating the model either completely at each retraining step or not at all. We excluded hyperparameter tuning due to its high cost and marginal expected benefit. We treated $r=1$ (or $r=7$ for daily data) as the accuracy benchmark and $r=T$ as the no-retraining baseline, with intermediate values representing periodic retraining strategies.
3.4 Evaluation metrics
3.4.1 Performance metrics
Evaluating point forecast accuracy in time series is a debated topic, with no consensus on the best metric (Hewamalage et al., 2023). In our context, characterized by intermittent data, metrics based on absolute or percentage errors are suboptimal (Kolassa, 2020), and due to varying scales across series, a scaled accuracy metric should be preferred. Hence, we adopted the Root Mean Squared Scaled Error (RMSSE) (Hyndman and Koehler, 2006),
$\text{RMSSE}=\sqrt{\frac{\frac{1}{h}\sum_{t=n+1}^{n+h}(y_{t}-\hat{y}_{t})^{2}}{\frac{1}{n-s}\sum_{t=s+1}^{n}(y_{t}-y_{t-s})^{2}}}\mbox{.}$ (3)
which compares the model’s squared error to that of a seasonal naive forecast, and it was the official accuracy metric in the M5 competition (with $s=1$) (Makridakis et al., 2022c). Lower values indicate better performance.
To assess probabilistic accuracy, we evaluated the quantiles produced using a scaled version of the Multi-Quantile Loss. The Scaled Quantile Loss, also known as (Scaled) Pinball Loss, and the Scaled Multi-Quantile Loss are defined as:
$\text{SQL}(\alpha)=\frac{\frac{1}{h}\sum_{t=n+1}^{n+h}\Big(\alpha\cdot(y_{t}-\hat{q}_{t}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t}^{(\alpha)}\leq y_{t}}+(1-\alpha)\cdot(\hat{q}_{t}^{(\alpha)}-y_{t})\cdot\mathbb{I}_{\hat{q}_{t}^{(\alpha)}>y_{t}}\Big)}{\frac{1}{n-s}\sum_{t=s+1}^{n}|y_{t}-y_{t-s}|}\mbox{,}$ (4)
$\text{SMQL}=\frac{1}{|\mathcal{Q}|}\sum_{\alpha\in\mathcal{Q}}\text{SQL}(\alpha)\mbox{,}$ (5)
These proper scoring rules assess forecast distribution accuracy (Kolassa, 2016). Moreover, the SMQL was the official metric of the M5 Uncertainty competition (weighted, and with $s=1$) (Makridakis et al., 2022a).
We considered a total of 14 different quantiles of probability levels $\alpha\in$ {0.005, 0.025, 0.050, 0.100, 0.150, 0.200, 0.025, 0.750, 0.800, 0.850, 0.900, 0.950, 0.975, 0.995}. Therefore, we analyzed the 50% 60%, 70%, 80%, 90%, 95%, and 99% central prediction intervals. The 50% and 60% central prediction intervals provide a good description of the center of the forecast distribution, while the 90%, 95%, and 99% give more information about its tails. These quantiles provide sufficient information about the uncertainty of the forecasts and allow for the effective description of the whole distribution. To ensure reliable quantile estimates, conformal prediction intervals were computed on a validation set at least twice the forecast horizon, which constrained the number of series used  [^2] We used a validation set four times larger than the forecast horizon for the M4 and M5 datasets, and twice the forecast horizon for the VN1 dataset. This limited the number of time series from 4,227 to 2,557 for the M4 dataset, and from 30,490 to 28,298 for the M5 dataset. .
We also normalized each evaluation metric relative to the baseline retraining scenario, defined by the dataset frequency, to enable consistent comparison across models and retraining periods. To statistically validate our findings, we applied the Friedman-Nemenyi test (Demšar, 2006).
3.4.2 Stability metrics
The stability evaluation of point predictive models is a crucial topic in time series forecasting: very few metrics are available to capture models’ vertical stability, and there is no consensus in the literature on what metric to use (Godahewa et al., 2025). Van Belle et al. (2023) proposed the symmetric Mean Absolute Percentage Change (sMAPC), which measures the change of one to $h$-step ahead forecasts obtained by two consecutive forecast origins, providing a measurement of up to which extent the forecasts generated at the first origin are unstable compared to those generated at the second origin.
$\text{sMAPC}=\frac{200}{h-1}\sum_{t=n+1}^{n+h-1}\frac{\lvert\hat{y}_{t,n}-\hat{y}_{t,n-1}\rvert}{\lvert\hat{y}_{t,n}\rvert-\lvert\hat{y}_{t,n-1}\rvert}\mbox{.}$ (6)
Here, $\hat{y}_{t,n}$ and $\hat{y}_{t,n-1}$ are the forecasts generated for period $t$ with origins $n$ and n1n-1 respectively. The instability across different pairs of consecutive forecasting origins can be obtained by a simple average of the sMAPC values. Lower values imply more stable predictions.
Beyond point forecasts, our study also emphasizes the probabilistic stability of the models under different retraining strategies. Since most machine learning and deep learning models do not natively produce full predictive distributions, we employed Conformal Inference to construct prediction intervals around point forecasts. Its advantages (validity guarantees, minimal assumptions, computational simplicity, and effectiveness with limited data) make it especially suitable for our global forecasting setup, where different models are compared across heterogeneous datasets and varying retraining scenarios  [^3] One may prefer, instead, to train models using quantile losses to directly yield quantile forecasts, or to use inherently probabilistic models. These are also valid ways of modeling uncertainty, as demonstrated by the M5 Uncertainty competition (Makridakis et al., 2022a; Lainder and Wolfinger, 2022). However, comparing different probabilistic modeling frameworks is out of the scope of our analysis, and it may be an area of future research. . Because no metric exists to evaluate the vertical stability (or instability) of the resulting probabilistic quantile forecasts, we proposed a new measure, the Multi-Quantile Change (MQC), to comprehensively assess the stability of the probabilistic predictions. We defined the Quantile Change and the Multi-Quantile Change as:
$\text{QC}(\alpha)=\frac{1}{h-1}\sum_{t=n+1}^{n+h-1}\Big(\alpha\cdot(\hat{q}_{t,n-1}^{(\alpha)}-\hat{q}_{t,n}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t,n}^{(\alpha)}\leq\hat{q}_{t,n-1}^{(\alpha)}}+(1-\alpha)\cdot(\hat{q}_{t,n}^{(\alpha)}-\hat{q}_{t,n-1}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t,n}^{(\alpha)}>\hat{q}_{t,n-1}^{(\alpha)}}\Big)\mbox{,}$ (7)
$\text{MQC}=\frac{1}{|\mathcal{Q}|}\sum_{\alpha\in\mathcal{Q}}\text{QC}(\alpha)\mbox{,}$ (8)
where $\mathcal{Q}$ is the set of possible quantile levels $\alpha$. The Quantile Change (QC) metric is a measure of the change in forecasted quantiles, $\hat{q}_{t}^{(\alpha)}$, between two consecutive forecast origins. Given two forecast origins, $n$ and n1n-1, QC quantifies the average adjustment in the predicted quantiles across the forecast horizon $h$. The formula mirrors the Quantile Loss but replaces the actual observations $y_{t}$, with previous quantile forecasts $\hat{q}_{t,n-1}^{(\alpha)}$, thereby shifting the focus from forecast accuracy to forecast stability. The intuition is straightforward: if the forecasts are stable across time (i.e., $\hat{q}_{t,n}^{(\alpha)}\approx\hat{q}_{t,n-1}^{(\alpha)}$), then QC will be small. Large QC values indicate that the model updates its distribution significantly with each new observation, suggesting a lack of stability in its probabilistic output. By aggregating this across a set of quantile levels $\mathcal{Q}$, the Multi-Quantile Change (MQC) captures overall instability, providing a comprehensive metric for probabilistic stability.
One drawback of the MQC is that it is a scale-dependent measure of instability. To solve this issue, we also propose a scaled version of the MQC, the Scaled Multi-Quantile Change (SMQC), defined as
$\text{SQC}(\alpha)=\frac{\frac{1}{h-1}\sum_{t=n+1}^{n+h-1}\Big(\alpha\cdot(\hat{q}_{t,n-1}^{(\alpha)}-\hat{q}_{t,n}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t,n}^{(\alpha)}\leq\hat{q}_{t,n-1}^{(\alpha)}}+(1-\alpha)\cdot(\hat{q}_{t,n}^{(\alpha)}-\hat{q}_{t,n-1}^{(\alpha)})\cdot\mathbb{I}_{\hat{q}_{t,n}^{(\alpha)}>\hat{q}_{t,n-1}^{(\alpha)}}\Big)}{\frac{1}{n-s}\sum_{t=s+1}^{n}\left|y_{t}-y_{t-s}\right|}\mbox{,}$ (9)
$\text{SMQC}=\frac{1}{|\mathcal{Q}|}\sum_{\alpha\in\mathcal{Q}}\text{SQC}(\alpha)\mbox{,}$ (10)
where we adopted the in-sample one-step-ahead seasonal naive forecast as a scaling factor, to enable the effective application of this metric to forecasting models by ensuring that the results are not biased by the scale of individual time series.
To further contextualize the proposed metric, it is important to compare it with existing distributional distance metrics such as the Kullback–Leibler ($KL$) divergence and the Wasserstein distance ($W_{2}$). KL divergence measures the relative entropy between two probability distributions and is particularly sensitive to differences in the tails (Kullback and Leibler, 1951). However, it requires access to full predictive densities, which can limit its use in forecasting applications where outputs are often expressed only as quantiles. Wasserstein distance ($W_{2}$), by contrast, quantifies the minimal cost of transforming one distribution into another and effectively captures both location and shape differences, but it usually relies on full distributions (although it can be approximated through discretization) (Peyré and Cuturi, 2019). The MQC metric addresses a complementary objective: it evaluates the stability of probabilistic forecasts directly at the quantile level, without the need for full predictive densities. Since it is based solely on a finite set of quantiles rather than the full predictive distribution, MQC does not capture all aspects of the distributional change, potentially missing more subtle shifts in shape, skewness, or variance that occur between those quantiles. However, in practical forecasting applications, especially in retail and supply chain domains where decisions are often made based on specific quantile levels (e.g., high quantiles for safety stock), this limitation is mitigated by the fact that the selected quantiles are usually those of greatest operational interest. Importantly, MQC is both model-agnostic and distribution-free, meaning that it can be applied consistently across different forecasting approaches without requiring strong assumptions about the underlying predictive distributions. Moreover, MQC is computationally lightweight compared to the Wasserstein distance or the KL divergence, which makes it suitable for large-scale applications involving thousands of time series.
4 Results and discussion
This section discusses the empirical findings of our study, highlighting the interplay between accuracy and stability across different retraining strategies.
Figure 1: RMSSE and sMAPC results for each global method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M4 and M5 datasets, and $r=1$ for the VN1 dataset.
Figure 1 shows the point forecast accuracy (on the left) and the point forecast stability (on the right) of each model  [^4] For the long computing time required, the Random Forest model has not been trained on the M5 dataset, while the LSTM model has not been adopted on both the M4 and M5 datasets. Moreover, the CatBoost method has not been used on the M4 dataset since too few categorical variables are available for this dataset. along the different retraining scenarios for the M4, M5, and VN1 datasets. To allow comparisons, we display the results in relative terms with respect to the baseline scenarios, that is $r=7$ for M4 and M5, and $r=1$ for VN1. In terms of forecast stability, the models’ behavior is similar across the different datasets. As expected, the sMAPC profiles are non-increasing functions of the retraining scenario. The stability remains practically the same, or even improves, as the retraining period increases. Indeed, moving from a high to a low retraining scenario, the stability of most global models improves compared to the baseline. In particular, for some models (i.e. CatBoost, LSTM, and TCN), the stability improvement is strong and consistent, reaching very low levels. These results imply that less frequent retraining may have a huge impact on the point forecast stability of global models. This can be explained by the fact that, when a forecasting origin is updated, and the model is retrained on the new data, the model’s predictions may differ consistently (even using the cross-learning approach) from those produced before the update. Therefore, less retraining is often synonymous with more stability. The statistical tests (see the Supplementary material) confirm the above results, showing that most periodic retraining scenarios are statistically different to the continuous retraining strategy in terms of point forecast stability at the 5% level. Similar considerations arise for the point forecast accuracy: the RMSSE profiles of the global models remain remarkably stable across retraining scenarios, sometimes displaying even modest improvements in accuracy when retrained less frequently (as is the case of deep learning models for the M4 dataset). In particular, under low periodic retraining schemes, the performance of most global models is statistically indistinguishable from the baseline configuration, and even under extreme settings (such as the no-retraining scenario) any degradation in accuracy remains below 5%. Overall, these findings show that global models can preserve, and in some cases even enhance, their forecasting accuracy over extended retraining intervals. Hence, frequent retraining may introduce unnecessary variability without delivering tangible gains in both predictive performance and forecast stability.
Figure 2: SMQL and SMQC results for each method and retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$ for the M4 and M5 datasets, and $r=1$ for the VN1 dataset.
In a similar fashion, Figure 2 summarizes the relative performance (on the left) and stability (on the right) from a probabilistic forecasting perspective. In this context, we observe that, for daily data, the stability paths are constant across almost every retraining scenario, meaning that less frequent updates neither harm nor improve the probabilistic forecasting stability of most global models  [^5] There are a few spikes in the SMQC profiles of some models. These can be due to different factors (e.g., inappropriate hyperparameter values). Nevertheless, the profiles are stable overall. . For weekly data, instead, we observe some differences among the models. Some of them show constant improvements in stability as the retraining scenario increases, while others do not exhibit particular enhancements or deteriorations in most periodic retraining intervals. When shifting the focus from point to probabilistic forecasting, the impact of retraining frequency on accuracy becomes more pronounced. Unlike point forecasts, probabilistic accuracy is more sensitive to longer retraining intervals. In several cases, reducing the retraining frequency leads to a gradual deterioration in probabilistic performance, indicating that the estimation of predictive uncertainty is more affected by outdated model parameters than point predictions. Nevertheless, these effects remain limited for low periodic retraining scenarios, where differences in probabilistic accuracy are typically negligible (below 2%). Importantly, this sensitivity is not universal: some models exhibit clear accuracy losses, while others maintain stable probabilistic performance or even improve. Overall, these findings suggest that accuracy uncertainty estimates benefit more from regular updates, yet do not strictly require very frequent retraining to remain reliable.
Figure 3: RMSSE, sMAPC, SMQL, and SMQC results for ETS method on the M5 dataset for each retrain scenario combination in relative terms with respect to the baseline scenario, $r=7$.
To further contextualize the stability findings for global models, we also evaluated two classical local forecasting approaches: automatic ARIMA and automatic ETS  [^6] For the long computing time required, the ARIMA model has not been trained on the M5 dataset. . Both local models exhibit highly consistent behaviors across datasets and retraining scenarios (see Supplementary material). As an illustrative example, Figure 3 reports the results for ETS on the M5 dataset. The accuracy profiles for both point and probabilistic forecasting, measured by RMSSE and SMQL respectively, are strictly increasing functions of the retraining period, indicating a systematic deterioration in performance as retraining becomes less frequent. Conversely, as expected, the stability profiles, captured by sMAPC and SMQC, are strictly decreasing, showing that less frequent retraining leads to progressively more stable forecasts. This monotonic trade-off highlights an intrinsic property of local models: stability improvements are usually obtained at the cost of accuracy. In contrast, global models display markedly different dynamics, often maintaining both accuracy and stability over a wide range of retraining periods, and, in some cases, even improving along one dimension without compromising the other. This comparison suggests that global forecasting models are inherently more robust to reduced retraining, benefiting from cross-series learning that mitigates the sensitivity to individual time series updates, a property that local models, by construction, cannot exploit.
Overall, the results from Figures 1 and 2, supported by the statistical tests, indicate that reducing the retraining frequency of global models does not negatively impact (and may even enhance) forecast stability, for both point and probabilistic forecasts. These results indicate that the accuracy–stability trade-off in global forecasting models is substantially milder than in local approaches. When considered alongside the findings of Zanotti (2025) on energy consumption, this provides strong evidence against the practice of frequent retraining in global forecasting models. Lower retraining scenarios preserve both accuracy and stability, while offering substantial savings in computational resources and energy consumption of forecasting systems.
Figure 4: The optimal retraining frequency identified for each forecasting objective, and for each series in the datasets considering all the models.
Finally, to jointly assess how accuracy- and stability-driven objectives translate into different retraining policies, Figure 4 provides a distributional view of the optimal retraining frequency across time series, in both point and probabilistic forecasting settings. Overall, the results indicate that continuous retraining is not required to achieve either accurate or stable forecasts when using global models. Across both point and probabilistic objectives, the retraining frequency can be safely reduced in favor of a periodic strategy, with stability-based criteria generally favoring slightly less frequent updates, in some cases even no retraining. However, compared to point forecasting, probabilistic objectives tend to favor slightly more frequent retraining, suggesting that accurate and stable uncertainty estimation requires a more regular model re-estimation. In practical terms, for daily data, a retraining interval of approximately 21–30 days emerges as a reasonable compromise across objectives, while for weekly data, retraining can be postponed usually up to 8 weeks without compromising performance. These retraining schedules strike a balance between the need to update models in response to potential distributional changes and the desire to limit unnecessary updates. Importantly, adopting such periodic retraining strategies can reduce the computational cost of forecasting by up to 75% Zanotti (2025), leading to more efficient and sustainable forecasting systems without sacrificing accuracy and also improving forecast stability.
5 Conclusions
This study investigates the role of retraining frequency in global forecasting models, with a particular focus on its implications for both forecast accuracy and forecast stability in point and probabilistic settings. Using three large-scale benchmark datasets (M4, M5, and VN1), we provided an extensive empirical analysis across multiple machine learning and deep learning architectures, complemented by a comparison with classical local forecasting methods. These results contribute to a more nuanced understanding of retraining as a design choice in large-scale forecasting systems, rather than as a default operational requirement.
Across all datasets and modeling families, we find that reducing the retraining frequency does not harm and often improves forecast stability. This result holds for both point and probabilistic forecasts, as measured by the proposed Scaled Multi-Quantile Change (SMQC) metric. In particular, stability profiles are typically flat or monotonically improving as retraining becomes less frequent, indicating that frequent retraining often introduces unnecessary variability in forecasts without delivering tangible benefits. These results are statistically validated via statistical tests and indicate that continuous retraining is often not the optimal strategy.
While point forecast accuracy is largely unaffected by reduced retraining, remaining stable or even improving in many cases, probabilistic accuracy shows a slightly higher sensitivity to retraining frequency. Less frequent updates may lead to modest degradations in distributional accuracy, although these effects are generally negligible for low periodic retraining scenarios and remain limited even under no-retraining regimes. Contrary to the common view of a strict trade-off between forecasting accuracy and stability, our results indicate that, in the context of global forecasting models, this tension is largely attenuated. Moreover, in many cases, reducing the retraining frequency does not require sacrificing predictive performance and may even contribute to both more stable and more accurate forecasts. This behavior contrasts with what is often observed in local modeling settings, where model re-estimation can lead to pronounced fluctuations in both accuracy and forecast revisions. For global models, the shared learning across series appears to act as a regularizing mechanism, making them less sensitive to frequent updates and allowing lower retraining frequencies to jointly improve stability while preserving, or even enhancing, also accuracy.
From a practical standpoint, these findings have several implications. First, businesses that rely on large-scale forecasting systems, such as those in retail, energy, and supply chain planning can adopt lower retraining strategies not only to reduce computational resources and the associated costs of forecasting systems, but also to increase operational stability, without compromising on forecast accuracy. Indeed, our results strongly support low, periodic retraining strategies (i.e. every 30 days or every 8 weeks) in global forecasting systems. Finally, these results promote a shift in emphasis from pure accuracy to stability-aware model selection and evaluation, especially in those contexts that are more sensitive to forecast reviews.
Nevertheless, some limitations of the present study should be acknowledged. While the analysis covers a broad range of global models and three state-of-the-art forecasting datasets, exploring the effects of retraining newer methods and different datasets may be a concrete area for future research. Moreover, this work focuses on the stability of global models. A deeper analysis of local approaches could therefore represent a valuable direction for future research. Additionally, the study focuses solely on full retraining schemes and does not explore incremental or online learning strategies, which may offer a different trade-off between accuracy and stability. Future research could extend this work by incorporating such adaptive methods and introducing business-oriented stability metrics that link forecast stability directly to operational cost or risk. Future research should also investigate the interplay between retraining frequency and hyperparameter optimization. Different models may exhibit heterogeneous responses to retraining strategies, depending on their hyperparameter configurations, and a deeper understanding of these interactions could enable the design of even more efficient and robust forecasting strategies. Furthermore, this study employed a single framework for generating probabilistic forecasts, namely Conformal Inference. Although conformalized residuals provide a valid means for quantifying uncertainty, future research could address this limitation by studying probabilistic stability in native probabilistic models, thereby enhancing the robustness and generalizability of the findings. Finally, further theoretical development of stability measures, particularly for probabilistic forecasts, may also contribute to a deeper understanding of how models behave across time and under varying data regimes.
In summary, this work provides strong empirical evidence that frequent retraining is neither necessary nor desirable in many forecasting applications. By explicitly accounting for stability alongside accuracy, forecasting systems can be designed to be more robust and computationally sustainable.
Supplementary material
Model Learning rate Iterations Feature fraction
Random Forest 0.03 1200 0.85
XGBoost 0.03 1200 0.85
LightGBM 0.03 1200 0.85
CatBoost 0.03 1200
Table 2: Main hyperparameter settings for machine learning forecasting models.
Model Learning rate Epochs Scaler
MLP 0.001 1000 Robust
LSTM 0.001 1000 Robust
TCN 0.001 1000 Robust
NBEATS 0.001 1000 Robust
NHITS 0.001 1000 Robust
Table 3: Main hyperparameter settings for deep learning forecasting models.
Method 7 14 21 30 60 90 120 150 180 364
ETS 1.606 1.760 1.925 2.035 2.510 3.041 3.341 3.704 4.113 5.449
ARIMA 1.621 1.776 1.944 2.056 2.545 3.094 3.356 3.804 4.276 5.674
LR 1.578 1.578 1.578 1.578 1.580 1.578 1.581 1.579 1.581 1.584
RF 1.715 1.713 1.713 1.710 1.707 1.699 1.700 1.701 1.701 1.706
XGBoost 18.125 18.115 18.140 18.142 18.144 18.116 18.130 18.026 18.088 18.112
LGBM 15.467 15.452 15.450 15.435 15.413 15.393 15.386 15.282 15.261 15.034
MLP 1.705 1.611 1.676 1.562 1.533 1.546 1.526 1.536 1.529 1.526
TCN 2.280 2.043 1.997 1.815 1.727 1.733 1.699 1.758 1.761 1.730
NBEATSx 1.625 1.521 1.494 1.475 1.468 1.471 1.469 1.473 1.475 1.498
NHITS 1.629 1.548 1.522 1.506 1.493 1.494 1.496 1.496 1.499 1.503
Table 4: M4 RMSSE values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
ETS 0.014 0.010 0.008 0.007 0.005 0.004 0.003 0.003 0.002 0.000
ARIMA 0.014 0.010 0.008 0.007 0.005 0.004 0.003 0.003 0.002 0.000
LR 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013
RF 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013
XGBoost 0.052 0.052 0.052 0.052 0.052 0.052 0.052 0.052 0.052 0.052
LGBM 0.123 0.123 0.123 0.123 0.123 0.122 0.122 0.122 0.122 0.121
MLP 0.017 0.016 0.017 0.016 0.016 0.016 0.016 0.016 0.016 0.016
TCN 0.021 0.017 0.016 0.015 0.015 0.015 0.015 0.015 0.014 0.015
NBEATSx 0.015 0.015 0.015 0.015 0.015 0.015 0.015 0.015 0.015 0.016
NHITS 0.016 0.015 0.015 0.015 0.015 0.015 0.015 0.015 0.015 0.015
Table 5: M4 sMAPC values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
ETS 0.419 0.467 0.513 0.549 0.684 0.864 0.948 1.078 1.230 1.563
ARIMA 0.428 0.475 0.524 0.560 0.701 0.881 0.955 1.111 1.276 1.655
LR 0.394 0.394 0.393 0.395 0.397 0.398 0.397 0.397 0.398 0.400
RF 0.437 0.438 0.437 0.439 0.440 0.440 0.439 0.439 0.440 0.443
XGBoost 2.279 2.288 2.293 2.309 2.337 2.365 2.394 2.405 2.447 2.636
LGBM 2.455 2.473 2.480 2.511 2.564 2.624 2.680 2.707 2.793 3.096
MLP 0.501 0.465 0.470 0.443 0.437 0.439 0.433 0.436 0.437 0.429
TCN 0.652 0.583 0.577 0.531 0.508 0.532 0.506 0.528 0.521 0.507
NBEATSx 0.466 0.426 0.415 0.410 0.409 0.409 0.410 0.407 0.409 0.421
NHITS 0.466 0.435 0.423 0.419 0.414 0.414 0.414 0.412 0.414 0.415
Table 6: M4 SMQL values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
ETS 0.278 0.300 0.326 0.351 0.396 0.435 0.449 0.469 0.493 0.603
ARIMA 0.279 0.301 0.328 0.347 0.397 0.439 0.451 0.482 0.496 0.593
LR 0.262 0.262 0.262 0.261 0.263 0.264 0.266 0.266 0.268 0.276
RF 0.351 0.351 0.351 0.352 0.352 0.353 0.354 0.353 0.353 0.357
XGBoost 2.156 2.150 2.154 2.153 2.149 2.142 2.134 2.126 2.130 2.119
LGBM 1.935 1.927 1.928 1.915 1.906 1.904 1.890 1.864 1.861 1.795
MLP 0.371 0.359 0.369 0.346 0.349 0.353 0.354 0.357 0.350 0.360
TCN 0.506 0.397 0.375 0.352 0.323 0.345 0.343 0.332 0.320 0.331
NBEATSx 0.338 0.330 0.330 0.330 0.331 0.333 0.341 0.337 0.341 0.354
NHITS 0.344 0.332 0.329 0.328 0.326 0.326 0.330 0.328 0.331 0.330
Table 7: M4 SMQC values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
ETS 0.747 0.751 0.756 0.759 0.769 0.776 0.778 0.774 0.784 0.782
LR 0.759 0.759 0.759 0.759 0.758 0.758 0.758 0.758 0.759 0.758
XGBoost 0.737 0.737 0.737 0.737 0.737 0.737 0.737 0.737 0.736 0.735
LGBM 0.753 0.753 0.753 0.753 0.753 0.753 0.753 0.753 0.753 0.753
CatBoost 0.917 0.929 0.924 0.928 0.921 0.936 0.925 0.930 0.934 0.926
MLP 0.803 0.800 0.800 0.801 0.801 0.802 0.799 0.801 0.801 0.800
TCN 0.844 0.844 0.844 0.844 0.844 0.844 0.844 0.843 0.843 0.841
NBEATSx 0.797 0.796 0.796 0.807 0.797 0.797 0.797 0.797 0.797 0.796
NHITS 0.811 0.808 0.817 0.808 0.808 0.806 0.807 0.807 0.807 0.805
Table 8: M5 RMSSE values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
ETS 0.104 0.070 0.054 0.045 0.026 0.018 0.012 0.013 0.007 0.000
LR 0.188 0.188 0.188 0.188 0.189 0.190 0.190 0.191 0.192 0.196
XGBoost 0.075 0.074 0.073 0.073 0.072 0.073 0.072 0.072 0.073 0.073
LGBM 0.054 0.053 0.053 0.053 0.052 0.052 0.052 0.052 0.052 0.052
CatBoost 0.120 0.045 0.060 0.039 0.031 0.022 0.021 0.020 0.020 0.023
MLP 0.499 0.511 0.505 0.519 0.512 0.524 0.510 0.525 0.475 0.552
TCN 0.497 0.208 0.174 0.086 0.055 0.043 0.031 0.017 0.017 0.003
NBEATSx 0.547 0.540 0.527 0.560 0.528 0.535 0.522 0.536 0.546 0.424
NHITS 0.566 0.547 0.557 0.533 0.566 0.536 0.532 0.548 0.511 0.565
Table 9: M5 sMAPC values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
ETS 0.249 0.251 0.254 0.256 0.260 0.263 0.264 0.264 0.267 0.288
LR 0.267 0.268 0.269 0.268 0.270 0.271 0.272 0.271 0.272 0.273
XGBoost 0.245 0.246 0.247 0.247 0.249 0.250 0.252 0.251 0.251 0.253
LGBM 0.246 0.247 0.247 0.247 0.248 0.249 0.250 0.249 0.250 0.251
CatBoost 0.277 0.268 0.278 0.272 0.273 0.269 0.265 0.271 0.268 0.267
MLP 0.263 0.264 0.265 0.266 0.269 0.271 0.271 0.272 0.275 0.276
TCN 0.270 0.271 0.272 0.273 0.276 0.278 0.279 0.279 0.281 0.282
NBEATSx 0.262 0.263 0.264 0.268 0.268 0.270 0.271 0.271 0.274 0.275
NHITS 0.266 0.266 0.273 0.268 0.271 0.273 0.274 0.273 0.276 0.278
Table 10: M5 SMQL values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
ETS 0.105 0.105 0.105 0.104 0.103 0.102 0.102 0.101 0.103 0.102
LR 0.110 0.110 0.110 0.111 0.111 0.111 0.112 0.112 0.112 0.113
XGBoost 0.100 0.099 0.099 0.100 0.100 0.100 0.099 0.100 0.099 0.098
LGBM 0.118 0.117 0.117 0.119 0.118 0.118 0.118 0.119 0.118 0.118
CatBoost 0.156 0.157 0.154 0.157 0.157 0.157 0.157 0.156 0.156 0.155
MLP 0.083 0.083 0.083 0.083 0.084 0.084 0.085 0.085 0.083 0.085
TCN 0.083 0.083 0.083 0.082 0.082 0.082 0.082 0.082 0.080 0.080
NBEATSx 0.084 0.084 0.084 0.088 0.085 0.085 0.085 0.086 0.088 0.105
NHITS 0.087 0.085 0.101 0.086 0.087 0.088 0.088 0.091 0.088 0.089
Table 11: M5 SMQC values for each method and retrain scenario combination.
Method 1 2 3 4 6 8 10 13 26 52
ETS 1.886 1.910 2.033 1.967 2.032 2.050 2.146 2.398 2.113 2.036
ARIMA 1.935 1.944 1.969 1.996 1.897 1.789 2.010 1.898 1.821 1.830
LR 6.886 6.873 6.925 6.886 6.892 6.888 6.870 6.947 7.037 7.077
RF 1.871 1.873 1.872 1.868 1.867 1.866 1.861 1.863 1.852 1.840
XGBoost 1.898 1.871 1.872 1.865 1.842 1.851 1.809 1.833 1.851 1.797
LGBM 3.540 3.520 3.471 3.282 3.414 3.732 3.415 3.337 3.357 3.543
CatBoost 5.631 5.570 5.694 5.025 4.846 5.359 4.418 4.876 5.212 5.343
MLP 1.554 1.483 1.477 1.472 1.472 1.475 1.480 1.480 1.467 1.455
LSTM 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960
TCN 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960
NBEATSx 1.704 1.696 1.661 1.561 1.519 1.525 1.544 1.462 1.455 1.456
NHITS 1.705 1.680 1.629 1.556 1.547 1.460 1.528 1.453 1.455 1.451
Table 12: VN1 RMSSE values for each method and retrain scenario combination.
Method 1 2 3 4 6 8 10 13 26 52
ETS 0.156 0.114 0.093 0.070 0.054 0.039 0.032 0.034 0.014 0.000
ARIMA 0.150 0.098 0.080 0.060 0.046 0.034 0.027 0.030 0.012 0.000
LR 0.086 0.084 0.085 0.083 0.083 0.086 0.086 0.085 0.084 0.083
RF 0.093 0.090 0.090 0.090 0.088 0.089 0.089 0.090 0.089 0.090
XGBoost 0.146 0.129 0.125 0.120 0.118 0.115 0.114 0.119 0.116 0.113
LGBM 0.121 0.110 0.111 0.106 0.104 0.105 0.105 0.104 0.104 0.105
CatBoost 0.312 0.197 0.185 0.152 0.134 0.139 0.115 0.117 0.101 0.095
MLP 0.457 0.468 0.467 0.475 0.442 0.483 0.437 0.492 0.468 0.451
LSTM 0.557 0.301 0.202 0.126 0.101 0.044 0.055 0.048 0.032 0.013
TCN 0.599 0.259 0.226 0.177 0.120 0.091 0.052 0.056 0.020 0.013
NBEATSx 0.594 0.592 0.594 0.576 0.576 0.581 0.572 0.580 0.571 0.549
NHITS 0.605 0.587 0.581 0.571 0.531 0.540 0.550 0.527 0.574 0.581
Table 13: VN1 sMAPC values for each method and retrain scenario combination.
Method 1 2 3 4 6 8 10 13 26 52
ETS 1.003 1.032 1.077 1.084 1.115 1.133 1.159 1.249 1.196 1.259
ARIMA 0.914 0.936 0.953 0.955 0.974 0.951 1.016 1.021 1.044 1.141
LR 4.270 4.292 4.412 4.378 4.451 4.294 4.295 4.454 4.358 4.669
RF 1.220 1.230 1.232 1.224 1.237 1.254 1.257 1.284 1.313 1.326
XGBoost 1.503 1.496 1.493 1.505 1.483 1.488 1.466 1.451 1.515 1.479
LGBM 2.318 2.292 2.230 2.157 2.230 2.317 2.262 2.362 2.220 2.268
CatBoost 4.203 4.122 4.292 3.970 3.735 4.028 3.549 3.798 3.863 4.192
MLP 0.806 0.800 0.796 0.795 0.803 0.805 0.807 0.819 0.824 0.847
LSTM 0.892 0.895 0.899 0.903 0.910 0.917 0.924 0.935 0.957 1.003
TCN 0.891 0.895 0.899 0.903 0.910 0.917 0.924 0.935 0.957 1.003
NBEATSx 0.991 0.995 0.966 0.897 0.861 0.866 0.893 0.825 0.839 0.875
NHITS 0.992 0.996 0.944 0.894 0.876 0.814 0.880 0.820 0.833 0.862
Table 14: VN1 SMQL values for each method and retrain scenario combination.
Method 1 2 3 4 6 8 10 13 26 52
ETS 0.533 0.460 0.428 0.375 0.347 0.323 0.298 0.296 0.228 0.151
ARIMA 0.463 0.396 0.345 0.304 0.291 0.260 0.255 0.235 0.182 0.132
LR 1.559 1.551 1.545 1.536 1.549 1.525 1.544 1.632 1.674 1.894
RF 0.729 0.732 0.735 0.737 0.735 0.738 0.743 0.734 0.731 0.734
XGBoost 0.629 0.598 0.600 0.584 0.585 0.575 0.556 0.613 0.599 0.589
LGBM 1.480 1.416 1.449 1.471 1.502 1.349 1.455 1.394 1.593 1.445
CatBoost 2.265 1.859 1.765 1.582 1.546 1.577 1.329 1.576 1.302 1.243
MLP 0.219 0.281 0.278 0.278 0.282 0.275 0.266 0.285 0.267 0.283
LSTM 0.242 0.238 0.237 0.235 0.233 0.229 0.227 0.224 0.214 0.187
TCN 0.241 0.238 0.237 0.235 0.233 0.229 0.227 0.224 0.214 0.187
NBEATSx 0.765 0.763 0.708 0.567 0.477 0.442 0.520 0.348 0.362 0.369
NHITS 0.755 0.779 0.660 0.555 0.506 0.333 0.507 0.348 0.343 0.365
Table 15: VN1 SMQC values for each method and retrain scenario combination.
Figure 5: RMSSE and sMAPC results for each local method and retrain scenario combination in relative terms with respect to the baseline scenario, r = 7 for the M4 and M5 datasets, and r = 1 for the VN1 dataset.
Figure 6: SMQL and SMQC results for each local method and retrain scenario combination in relative terms with respect to the baseline scenario, r = 7 for the M4 and M5 datasets, and r = 1 for the VN1 dataset.
Figure 7: M4 Friedman-Nemenyi test results based on RMSSE.
Figure 8: M4 Friedman-Nemenyi test results based on sMAPC.
Figure 9: M4 Friedman-Nemenyi test results based on SMQL.
Figure 10: M4 Friedman-Nemenyi test results based on SMQC.
Figure 11: M5 Friedman-Nemenyi test results based on RMSSE.
Figure 12: M5 Friedman-Nemenyi test results based on sMAPC.
Figure 13: M5 Friedman-Nemenyi test results based on SMQL.
Figure 14: M5 Friedman-Nemenyi test results based on SMQC.
Figure 15: VN1 Friedman-Nemenyi test results based on RMSSE.
Figure 16: VN1 Friedman-Nemenyi test results based on sMAPC.
Figure 17: VN1 Friedman-Nemenyi test results based on SMQL.
Figure 18: VN1 Friedman-Nemenyi test results based on SMQC.
Figure 19: M4 Friedman-Nemenyi test results for local models.
Figure 20: M5 Friedman-Nemenyi test results for local models.
Figure 21: VN1 Friedman-Nemenyi test results for local models.